OpenStudy (anonymous):

Explain, in complete sentences, how you would completely factor 20x2 – 28x – 48 by using the grouping method. How would you check your factors for accuracy?

OpenStudy (anonymous):

=4(5x^2-7x-12)=4(5x^2+5x-12x-12)=4[5x(x+1)-12(x+1)]=4(x+1)(5x-12)